PLEASE HELP ME FAST

Lane Company manufactures a single product and applies overhead cost to that product using standard direct labor-hours. The bRequired: 1. Compute the predetermined overhead rate for the year. Break the rate down into variable and fixed elements. 2. Peqamed. Compute the predetermined overhead rate for the year. Break the rate down into variable and fixed elements. Prepare a1. Compute the predetermined overhead rate for the year. Break the rate down into variable and fixed elements. 2. Prepare a s1. Compute the predetermined overhead rate for the year. Break the rate down into variable and fixed elements. 2. Prepare a s
Answer #1

Solution 1:

Computation of Predetermined overhead rate - Lane Company
Particulars Amount
Budgeted variable manufacturing overhead ($4.20*195000) $819,000.00
Budgeted fixed manufacturing overhead $1,599,000.00
Total Manufacturing overhead $2,418,000.00
Budgeted direct labor hours 195000
Predetermined overhead rate $12.40
Variable overhead rate (Per direct labor hour) $4.20
Fixed manufacturing overhead rate (per direct labor hour) $8.20

Solution 2a:

Standard Cost Card - Lane Company
Particulars Amount
Direct Material (4*$8.50) $34.00
Direct labor (1.5*$13.10) $19.65
Variable manufacturing overhead (1.5*$4.20) $6.30
Fixed manufacturing overhead (1.5*$8.20) $12.30
Standard cost per unit $72.25

Solution 3a:

Standard hours for actual production = 156000*1.50 = 234000 hours

Solution 3b:

Manufacturing Overhead
Particulars Debit Particulars Credit
To Cash (Variable manufacturing overhead) $633,750.00 By WIP (Applied overhead) (156000*1.50*$12.40) $2,901,600.00
To Cash (Fixed manufacturing overhead) $1,774,500.00
To overhead variance (Overapplied overhead) $493,350.00
Total $2,901,600.00 Total $2,901,600.00

solution 4:

Variable overhead actual rate = $633750 / 253500 = $2.50
Variable overhead rate variance = (SP - AP) *Actual Hours = ($4.20 - $2.50) * 253500 = $430950 Favorable
Variable overhead Efficiency variance = (Standard hours - Actual Hours) *SP = (156000*1.5 - 253500)*$4.20 = - $81,900 (Unfavorable)
Fixed overhead budget Variance = Budgeted Fixed Overhead - Actual Fixed overhead = $1,599000 - $1,774,500 = -$175,500 (unfavorable)
Fixed Overhead Volume Variance = ($8.20*156000*1.50) - $1,599,000 =$1,918,800 - 1,599,000 = $319,800 Favorable